# Complete moduli of Enriques surfaces in Horikawa’s model

### 06 Dec 2023 v1

@@ 6.1. Type III ADE surfaces
The only ADE surfaces needed in this paper are the ones that appear on the boundary of the KSBA compactification ${\overline{F}}_{(2,2,0)}$. They are described in detail in the last section of [AE22]. Most of them are easy: they are hypersurfaces in projective toric varieties in a way very similar to Horikawa’s construction in Section 2.1.
@@ 6.4. Anticanonical ADE surfaces with two commuting involutions
One could say that the ADE surfaces $Z^{\prime}\to W$ are obtained by folding the ADE surfaces $X\to Y$ by the symplectic involution $\iota_{{\rm Nik}}$, and $Z\to W$ are obtained from $X\to Y$ by folding by the nonsymplectic involution $\iota_{{\rm En}}$. The index-$1$ cover $\rho^{\prime}\colon Z^{\prime}\to W$ and the index $2$ cover $\rho\colon Z\to W$ are dual in the above sense.

### 28 Dec 2023 v2

@@ 6.1. Type III ADE surfaces
The only ADE surfaces needed in this paper are the ones that appear on the boundary of the KSBA compactification ${\overline{F}}_{(2,2,0)}$. They are described in detail in the last section of [AE22]. Most of them are easy: they are hypersurfaces in projective toric varieties in a way very similar to the construction in Section 2.1.
@@ 6.4. Anticanonical ADE surfaces with two commuting involutions
One could say that the ADE surfaces $Z^{\prime}\to W$ are obtained by folding the ADE surfaces $X\to Y$ by the symplectic involution $\iota_{{\rm Nik}}$, and $Z\to W$ are obtained from $X\to Y$ by folding by the nonsymplectic involution $\iota_{{\rm En}}$. The index-$1$ cover $\rho^{\prime}\colon Z^{\prime}\to W$ and the index $2$ cover $\rho\colon Z\to W$ are dual in a similar way to Remark 2.2.